OpenStudy (anonymous):

Let f be an odd function with domain that contains 0. Show that f(0)=0 Let g(x)=f(x)-f(-x) a. Show that g is an odd function b. If f is even show that g(x)=0 for all numbers x.

myininaya (myininaya):

well use g(x)=f(x)-f(-x) what is g(0)?

OpenStudy (anonymous):

not sure

myininaya (myininaya):

well if you replace x with 0 you have g(0)=f(0)-f(-0) ... we can find g(0) easily but then we need to somehow find that f(0)=0

myininaya (myininaya):

major hint: recall if f is odd f(-x)=-f(x)

myininaya (myininaya):

another hint: you want to know what happens when x is 0

myininaya (myininaya):

is there 3 questions here? or are you suppose to use the 2 questions beneath to show f(0)=0?

myininaya (myininaya):

you don't need those ...

OpenStudy (anonymous):

just two questions

myininaya (myininaya):

but anyways again replace x with 0 f(-x)=-f(x)

myininaya (myininaya):

I see 3 questions

myininaya (myininaya):

show f(0)=0 show g is odd f is even then show g=0 for all x

OpenStudy (anonymous):

ohh ok. Can you help me answer them

myininaya (myininaya):

trying to but you haven't used what I said to show f(0)=0

myininaya (myininaya):

you have f is odd so f(-x)=-f(x) and you want to know what happens at x=0 so plug in 0 for x

myininaya (myininaya):

anyways I guess I will move on to the next one... hint for "show g is odd" g(x)=f(x)-f(-x) to show g is odd you need to show g(-x)=-g(x) so plug in -x and show that

myininaya (myininaya):

hint for the last one: if f is even then g=0 for all x well recall g(x)=f(x)-f(-x) and since f is even then f(-x)=f(x)